OpenStudy (anonymous):

MEDALS!! Please help...If an investor pusrchased 50 shares of Pearson Education Stock at $85 per share, 90 shares at $105 per share, 120 shares at $110 per share and another 75 shares at $130 per share what is the mean cost per share?

OpenStudy (texaschic101):

50 at $85 = 50 x 85 = 4250 90 at $ 105 = 90 x 105 = 9450 120 at $ 110 = 120 x 105 = 12,600 75 at $130 = 75 x 130 = 9750 add the money spent : 4250 + 9450 + 12,600 + 9750 = 36,050 add the number of stocks you have bought : 50 + 90 + 120 + 75 = 335 now take the total money and divide by the total number of stocks.. 36,050/335 = 107.6119 rounds to $ 107.61(mean cost per share)

OpenStudy (anonymous):

Thank you Texaschic101

OpenStudy (texaschic101):

no problem...do you have any questions ?

OpenStudy (anonymous):

it makes sense seeing it written out. I just wasn't sure when I read the problem on how to get the final outcome. I think I was over thinking it. :-/

OpenStudy (texaschic101):

I do that sometimes. Make something harder then what it really is. Glad I could clear it up for you :)

OpenStudy (anonymous):

thanks again!!!